OpenStudy (anonymous):

Rationalize the denominator 1/(sqrt4)

OpenStudy (jdoe0001):

well.. what's \(\bf \sqrt{4} \quad ?\)

OpenStudy (anonymous):

\[1/\sqrt{4}\] well, \[\sqrt{4}=2\] which makes your equation \[1/\sqrt{4}=1/2 \]